OpenStudy (anonymous):

Riemann Integral/The Definite Integral help? @Calculus1

OpenStudy (anonymous):

I don't understand how to solve these by hand for example f(x)=sin (x) on the interval [0,pi] when n=8

OpenStudy (zarkon):

are you using right hand endpoints or something else?

OpenStudy (anonymous):

The exact question is Estimate the area under a single loop of the sine curve sin(x) by erecting on the region 8 circumscribed rectangles of equal width.

OpenStudy (zarkon):

well there are may ways to do this. you can really pick and points in the partition of your interval...I'm going to go with right hand endpoints so let \[x_i=i\cdot\Delta x\] where \[i=1,2,\ldots 8\] and \[\Delta x=\frac{\pi}{8}\] then compute \[\sum_{i=1}^{8}\sin(x_i)\Delta x\]

OpenStudy (anonymous):

perfect, thank you